OpenStudy (anonymous):

Find an explicit rule for the nth term of the sequence. 2, -8, 32, -128, ...

OpenStudy (anonymous):

my answer;

OpenStudy (anonymous):

so i was right, right?

OpenStudy (bradely):

2, -8, 32, -128, ... ratio= -8/2=-4 an= ar^n-1 =2(-4)^n-1

Parth (parthkohli):

Mondona, try plugging in \(n = 1, 2, 3\cdots\) in your answer. Does it produce the correct values?

OpenStudy (anonymous):

@bradely that was not correct

OpenStudy (anonymous):

that was my answer before

Parth (parthkohli):

- What is the first term? - What is the common ratio?